### v2

Abstract
We revisit the relationship between attention mechanisms and large kernel ConvNets in visual transformers and propose a new spatial attention named Large Kernel Convolutional Attention (LKCA). It simplifies the attention operation by replacing it with a single large kernel convolution. LKCA combines the advantages of convolutional neural networks and visual transformers, possessing a large receptive field, locality, and parameter sharing. We explained the superiority of LKCA from both convolution and attention perspectives, providing equivalent code implementations for each view. Experiments confirm that LKCA implemented from both the convolutional and attention perspectives exhibit equivalent performance. We extensively experimented with the LKCA variant of ViT in both classification and segmentation tasks. The experiments demonstrated that LKCA exhibits competitive performance in visual tasks. Our code has already been made publicly available at https://github.com/CatworldLee/LKCA.
Keywords:
ConvNet, Vision Transformer, Large Kernel CNN, Attention Mechanism
1 Introduction
Convolutional Neural Networks (CNNs) [1, 2] have been a common choice for visual encoders in modern visual models. However, the emergence of Vision Transformers (ViTs) [3] has expanded the options for the backbone network of models. ViTs have demonstrated excellent performance in various visual tasks [4, 5, 6, 7].
Figure 1: Two views to interpret LKCA. The Large Kernel Convolutional Attention (LKCA) can be understood From the perspective of convolution on the left and attention on the right. The effects of the two approaches are equivalent.
CNN and ViT each possess unique advantages. CNN is fundamentally translation equivariance and shifting window invariance [8], providing robustness to subtle changes in object positions and orientations in images, leading to significant success in various computer vision tasks. In contrast, ViT effectively captures global context information through its self-attention mechanism [9], enabling the modeling of long-range dependencies [10] and contextual relationships [11]. Despite the significant progress made by CNN and ViT over the years, they still have some drawbacks and challenges. CNN may face difficulties in capturing remote contextual information [12], as its primary focus is on local feature extraction, potentially limiting its ability to understand global context [13]. This limitation can result in performance degradation in tasks that require global context information, such as scene understanding, image captioning, or fine-grained recognition. On the other hand, ViT lacks the inherent spatial inductive bias that CNN possesses [14], which is specifically designed to leverage the spatial locality of images.
In order to integrate the strengths of both CNN and ViT, we propose a Large Kernel Convolutional Attention (LKCA) mechanism. The implementation of LKCA is straightforward, replacing the self-attention operation with a single large kernel convolution operation. LKCA not only inherits the translational invariance and spatial inductive bias advantages of ConvNets but also leverages the global modeling and long-range dependency capabilities of ViT. Deconstructing LKCA from the perspective of attention can be considered as a novel approach of spatial attention. The approach closest to our work is Visual Attention Network (VAN) [15], where they replace the multi-head self-attention mechanism with three types of small kernel convolutions: DW-Conv, DW-DConv, and 1$\times$1-Conv. In contrast, our method is simpler, directly employing one large kernel convolution [12] to replace MHSA.
We extensively validated the effectiveness of LKCA in two visual tasks: image classification and semantic segmentation. In image classification, we conducted experiments on classic datasets including CIFAR10, CIFAR100, SVHN, and Tiny-ImageNet, demonstrating the efficacy of LKCA across these datasets. Notably, in the case of Tiny-ImageNet, we performed comparative experiments against prominent variants of Vision Transformers ViT and achieved compelling and competitive results. For semantic segmentation, we selected the ADE20K dataset and conducted ablation experiments under different framework configurations. Through comparisons with multiple backbone networks, we verified the superior performance of LKCA in terms of both parameter efficiency and accuracy. This further underscores the excellence of LKCA in addressing semantic segmentation tasks.
2 Related Work
2.1 Visual Attentions
Attention mechanisms [16, 17, 18, 19, 20, 21] are inspired by the way humans naturally focus on important elements in their subconscious. Visual attention mechanisms [20, 21] can be viewed as a dynamic weighting process based on input image features. Visual attention methods can be broadly categorized into four types: channel attention [22, 23, 24, 25], spatial attention [26, 27, 28], temporal attention [29, 30], and branch attention [31, 32, 33], as well as other mixed combinations [34, 35, 36, 37, 38, 39, 40]. Non-Local Network [41] successfully employs self-attention for the first time to simulate non-local relationships in computer vision. CBAM [42] infers attention weights sequentially along both spatial and channel dimensions and can seamlessly integrate into any CNN architecture. ECA [22] module introduces a non-reductive local cross-channel interaction strategy, which efficiently enhances the attention module’s effectiveness through the effective implementation of 1D convolution. EncNet [23] significantly improves the spatial resolution of pixel-level semantic segmentation labels by introducing a context encoding module. OCNet [26] combines an interleaved sparse self-attention scheme with a traditional multi-scale context approach, effectively modeling dense relationships between pixels and providing richer contextual information. SAGAN [27] employs spatial attention to model long-range dependencies in image generation tasks. SCAN [29] achieves significantly superior performance in Video Person Re-Identification by refining both intra-sequence and inter-sequence feature representations through the introduction of temporal attention to video sequences. SKNet [31] enhances the model’s capability to capture objects at different scales by introducing branch attention to merge multiple branches with different kernel sizes.
2.2 Large Kernel ConvNets
Compared to small-kernel CNNs [1, 43], large-kernel CNNs [44, 45, 46, 47, 48, 49, 50, 12, 15] possess a broader effective receptive field and stronger shape preferences. However, older models like Inception [44, 45, 46] fell out of favor after VGG-Net [47], leading to a decline in the popularity of large-kernel models. A notable research outcome is the Global Convolutional Network (GCN) [48], which employs very large convolutional kernels of 1$\times$×K and K×$\times$1 to optimize semantic segmentation tasks. However, there are reports suggesting that large kernels may compromise ImageNet performance. On the other hand, the Local Relation Network (LRNet) [49] introduces a spatial aggregation operator (LRLayer) as a replacement for standard convolution, which can be seen as a dynamic convolution. ConvMixer [50], on the contrary, uses convolutional kernels as large as 9$\times$9 to replace the ”mixer” components of ViT or MLP. Inspired by ViT, RepLKNet [12] utilizes a large 31$\times$31 convolutional kernel, and its outstanding performance is primarily attributed to the effective receptive fields constructed through large kernels [65]. It has been demonstrated that RepLKNet better exploits shape information compared to traditional CNNs. The Visual Attention Network (VAN) [15] employs three types of convolutions—DW-Conv, DW-DConv, and 1$\times$1 Conv—to expand the receptive field of convolutions and integrate them into the attention mechanism.
2.3 Large ERF ViTs
The strength of ViTs lies in its large Effective Receptive Fields (ERFs). Transformer [51] was initially applied in the field of Natural Language Processing, and the first work to migrate its powerful representation capabilities to the Computer Vision domain was ViT [3]. Following the paradigm of ViT, a series of ViT variants [52, 53, 54, 55, 56, 57, 58, 59, 60, 61] have been proposed to enhance the performance of visual tasks. Visual Transformers excel at capturing long-range dependencies between patches, representing large receptive fields. However, ViT lacks the dependence on short-term relationships for 2D image local dependencies that CNNs inherently possess. Researchers have primarily focused on improving ViT’s modeling capability for local information. Transformer in Transformer (TNT) [52] enhances the feature representation capability of local patches through internal attention. Swin Transformer [53, 54] addresses local modeling challenges in adapting transformers from language to vision by introducing a hierarchical architecture with shift windowing. Twins [55] and CAT [56] alternately perform local and global attention layers layer-wise. Shuffle Transformer [57, 58] efficiently establishes inter-window connections through spatial shuffling with deep convolution and enhances short-term dependency modeling with neighboring window connections. RegionViT [59] introduces regions and local tokens with different patch sizes, enhancing local modeling through a region-to-local attention mechanism.
3 Method
3.1 Preliminary
3.1.1 Review of Large Kernel Convolution.
Ding et al. [12] revisit the paradigm of large kernel design in modern CNNs, drawing inspiration from the ViT. Ding et al. assert that employing a few large convolutional kernels instead of multiple smaller ones constitutes a more powerful design paradigm. Building upon this notion, the author introduces a novel CNN architecture named RepLKNet [12], featuring kernel sizes as large as 31$\times$31. In comparison to commonly used 3$\times$3 kernels, RepLKNet achieves performance results equivalent to or better than the Swin Transformer [53], with lower latency. The difference between small kernel convolution and large kernel convolution is shown in Figure 2
3.1.2 Comparision of LKC and MHSA.
Convolution possesses a natural advantage wherein parameters are shared based on geographical location. Specifically, identical parameters of the same convolutional kernel act on every position within the receptive field during the sliding window process. This introduces equivariance and translation invariance to the convolutional operation. However, in Vision Transformer, the attention scores generated by the self-attention mechanism do not explicitly exhibit positional invariance. Attention weights are determined by the feature vectors themselves and the introduced positional encodings. While this enhances the model’s fitting capability, it significantly increases the complexity of training. Introducing a mechanism for sharing parameters in the attention mechanism of ViT may prove beneficial in alleviating training challenges.
Figure 2: Illustration of the difference between small kernel convolution and large kernel convolution by constructing a 5x5 feature map, a 3x3 convolution kernel smaller than the feature map, and a 7x7 convolution kernel larger than the feature map. The distinction between small kernel convolution and large kernel convolution lies in the fact that, in small kernel convolution, all parameters of the kernel are involved in each correlation operation, while only a subset of feature map parameters participates in the computation. In the case of large kernel convolution, during each correlation operation, all parameters of the feature map are involved, and only a subset of the convolution kernel parameters participates in the computation.
3.2 Large Kernel Convolution Attention (LKCA)
3.2.1 Shared Weight Position Operation.
Firstly, we initialize a square matrix of size $N^{\frac{1}{2}}\times 2-1$ as learnable parameters, denoted as the shared parameter weights, where $N$ represents the number of patches, akin to the initialization of convolutional kernels. Subsequently, we initialize a window of size $N^{\frac{1}{2}}$ and traverse this window as it slides over the shared parameter weights matrix from the bottom-right corner to the top-right corner. At each sliding step, the matrix within the window is flattened row-wise into a row vector of size $(1,N)$ and stored. This process is iterated as the window slides, concatenating the row vectors obtained from each step. In total, the window slides $N$ times. Consequently, the concatenated matrix assumes a shape of $(N,N)$, aligning with the shape of the attention weight matrix. This concatenated matrix serves as the new attention score, undergoing matrix multiplication with the linearly mapped Value matrix. This step is referred to as the Shared Weight Matrix Attention. Listing 1.1 presents pseudocode for implementing LKCA from the attention perspective.
class Attention_ViT(nn.Module):
    def __init__(num_patches):
        super().__init__()
        n = num_patches
        H, W = n**0.5
        weight = nn.Parameter(torch.zeros(H*2-1, W*2-1))
    def forward(X):
        b, n, d = X.shape
        V = nn.Linear(X)
        for i in range(H):
            for j in range(W):
                start_row = H-1-i
                start_col = W-1-j
                window = weight[start_row:start_row+H, start_col:start_col+W]
                window_vector = window.flatten()
                attn_score[i*W+j, :] = window_vector
        Out = attn_score @ V
        return Out
Listing 1: Pseudocode from attention perspective
3.2.2 From the Perspectives of Convolution: Implementing LKCA.
Large Kernel Convolutional Attention (LKCA) can be expressed equivalently as the attention operation with shared parameters described above in terms of code and mathematical implementation. Listing 1.2 showcases pseudocode for implementing LKCA from the perspective of convolution. The convolutional kernel is a learnable weight matrix of size $N^{\frac{1}{2}}\times 2-1$, and the input feature map is a matrix $V$ of size $N^{\frac{1}{2}}$ with a channel count equal to the hidden dimension. After zero-padding of size $N^{\frac{1}{2}}-1$, the convolution operation is applied to the two matrices.
$\mathbf{LKCA(x)}\mathbf{=}\mathbf{Conv2d}(\mathbf{MLP(x)},\mathbf{LargeKernel})$
class Attention_CNN(nn.Module):
    def __init__(num_patches):
        super().__init__()
        n = num_patches
        H, W = n**0.5
        weight = nn.Parameter(torch.zeros(H*2-1, W*2-1))
    def forward(X):
        b, n, d = X.shape
        V= nn.Linear(X)
        V = rearrange(V, ’b (H W) d -> (b d) H W’)
        Out = F.conv2d(V, weight, zero_padding=(H-1, W-1))
        Out = rearrange(Out, ’(b d) H W -> b (H W) d’)
        return Out
Listing 2: Pseudocode from convolution perspective
3.2.3 Overall Architecture.
The incorporation of large-kernel convolution attention enhances the model’s ability to capture positional information. However, it simultaneously weakens the relationships between patches and within individual patches. Therefore, we intersperse the large-kernel convolution attention modules within the original ViT architecture while retaining the self-attention modules.
The model overview is represented by the following formulas. An image $x\in\mathbb{R}^{H\times W\times C}$, where $(H,W)$ is the resolution of the original image, and $C$ is the number of channels. A series of flattened 2D patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(P,P)$ is the resolution of each image block, and $N=\frac{HW}{P^{2}}$ is the number of generated blocks. The patches are linearly projected by weight matrix $\mathbf{E}$ to a $D$-dimensional space and added with position embedding $\mathbf{E}_{pos}$. Differing from the original ViT, we discard the use of the [class] token. Instead of extracting the [class] token from the final linear layer as classification information, we adopt the mean of the last layer’s feature vectors for classification. Additionally, we insert the Large Kernel Convolution Attention (LKCA) module before the Multi-Layer Perceptron (MLP) module and apply Layer Normalization (LN) before each module.
$\mathbf{z}_{0}$ $=\left[\mathbf{x}_{p}^{1}\mathbf{E};\mathbf{x}_{p}^{2}\mathbf{E};\cdots;\mathbf{x}_{p}^{N}\mathbf{E}\right]+\mathbf{E}_{pos},$ $\mathbf{E}\in\mathbb{R}^{\left(P^{2}\cdot C\right)\times D},\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}$ (1)
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{LKCA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L$ (2)
$\mathbf{z}_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}_{\ell}^{\prime}\right)\right)+\mathbf{z}_{\ell}^{\prime},$ $\ell=1\ldots L$ (3)
$\mathbf{y}$ $=\operatorname{LN}\left(\overline{\mathbf{z}_{L}}\right)$ (4)
4 Experiments
4.1 Image Classification
Experimental Setting. In our experimental setup, we conducted extensive image classification experiments on four prominent datasets: CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet. All four datasets are characterized by low resolution, with image sizes of 32 for CIFAR-10, CIFAR-100, and SVHN, and 64 for Tiny-ImageNet. The training regime was consistently maintained across all experiments, conducted on a single RTX 2080ti GPU. The training spanned 100 epochs, employing 4 data loading workers, a batch size of 128. For optimizer, we choose AdamW [62]. The default learning rate was set to 0.001, with a weight decay of 0.05. We utilized cosine annealing with a warm-up period of 10 epochs. During the training phase, we employed Stochastic Depth [63], Label Smoothing [45], and Random Erasing [64] with a probability of 0.25, a maximum erasing area of 0.4, and an aspect of erasing area of 0.3 as regularization strategies to enhance the robustness and generalization capabilities of our model. Regarding data augmentation, we employed various techniques. Specifically, we applied RandomHorizontalFlip, RandomCrop [65], Mixup [66], and Cutmix [67]. Additionally, we incorporated Autoaugmentation [68], utilizing the CIFAR10 Policy for CIFAR-10 and CIFAR-100, the SVHN Policy for SVHN, and the ImageNet Policy for Tiny-ImageNet.
Comparison of ViT and LKCA. Initially, we conducted extensive experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet to compare the performance of the original ViT with an LKCA-modified version of ViT having a similar order of magnitude in terms of parameters. The purpose of these experiments was to validate the effectiveness of LKCA across various parameter magnitudes and datasets. The parameter sizes of both the ViT and the LKCA-modified ViT varied across a range of magnitudes, spanning from small-scale models with 0.5 million parameters to large-scale models with 10 million parameters. Specifically, the parameter magnitudes were explored at levels of 0.5M, 1M, 2M, 4M, and 8M. It is important to note that the benchmark representing the model size across all variants is fixed at 2.69 million parameters, and it is independent of any subjective categorization as either large or small. This standardized model size serves as the reference point for assessing the performance of the different variants, ensuring a consistent comparison across the experiment.
Table 1: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-10 (Top-1 accuracy)
Model Acc # Params Flops
ViT 93.36 2.69M 174.25M
ViT-LKCA 94.11 (+0.75) 2.69M 170.81M
VIT/0.5M 87.83 0.48M 30.58M
ViT/0.5M-LKCA 90.94 (+3.11) 0.50M 31.34M
VIT/1M 91.16 1.06M 68.36M
ViT/1M-LKCA 92.49 (+1.33) 1.03M 64.82M
VIT/2M 92.32 1.96M 126.09M
ViT/2M-LKCA 93.55 (+1.23) 2.03M 128.66M
VIT/4M 93.64 3.95M 255.47M
ViT/4M-LKCA 94.71 (+1.07) 4.05M 257.43M
VIT/8M 94.46 7.85M 508.72M
ViT/8M-LKCA 94.88 (+0.42) 7.98M 506.00M
Table 1 presents a performance comparison of ViT and its LKCA variants on the CIFAR-10 image classification task, along with their performance across different orders of magnitude in terms of parameters. The table includes the Top-1 accuracy (Acc), the number of parameters (# Params), and Floating Point Operations (Flops) for each model. ViT and ViT-LKCA: The ViT model achieves a Top-1 accuracy of 93.36%, with 2.69M parameters and 174.25M Flops. The ViT-LKCA outperforms with a Top-1 accuracy of 94.11% (an improvement of 0.75 percentage points) and has 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model has a Top-1 accuracy of 87.83%, 0.48M parameters, and 30.58M Flops. The ViT/0.5M-LKCA, with LKCA improvement, significantly improves accuracy to 90.94% (an increase of 3.11 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): For ViT and ViT-LKCA models with 1M, 2M, and 4M parameters, LKCA consistently yields higher Top-1 accuracy. For instance, ViT/1M-LKCA achieves a 1.33 percentage point improvement, reaching an accuracy of 92.49%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance gap is narrower. ViT/8M-LKCA slightly improves Top-1 accuracy by 0.42 percentage points, reaching 94.88%.
Table 2: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-100 (Top-1 accuracy)
Model Acc # Params Flops
ViT 72.65 2.69M 174.25M
ViT-LKCA 76.50 (+3.85) 2.69M 170.81M
VIT/0.5M 64.68 0.48M 30.58M
ViT/0.5M-LKCA 69.35 (+4.67) 0.50M 31.34M
VIT/1M 70.86 1.06M 68.36M
ViT/1M-LKCA 73.40 (+2.54) 1.03M 64.82M
VIT/2M 71.17 1.96M 126.09M
ViT/2M-LKCA 75.94 (+4.77) 2.03M 128.66M
VIT/4M 72.82 3.95M 255.47M
ViT/4M-LKCA 76.95 (+4.13) 4.05M 257.43M
VIT/8M 73.39 7.85M 508.72M
ViT/8M-LKCA 78.57 (+5.18) 7.98M 506.00M
Table 2 illustrates Acomparison between ViT and LKCA in the context of the CIFAR-100. ViT and ViT-LKCA: The base ViT model achieves a Top-1 accuracy of 72.65% with 2.69M parameters and 174.25M Flops. In contrast, ViT-LKCA exhibits superior performance, attaining a Top-1 accuracy of 76.50% (an increase of 3.85 percentage points) with the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 64.68% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, featuring LKCA enhancements, significantly improves accuracy to 69.35% (a notable increase of 4.67 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): ViT and ViT-LKCA models with 1M, 2M, and 4M parameters consistently benefit from LKCA improvements. For instance, ViT/2M-LKCA achieves a noteworthy 4.77 percentage point improvement and reaches an accuracy of 75.94%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance difference is more subtle. ViT/8M-LKCA exhibits a huge improvement of 5.18 percentage points, reaching a Top-1 accuracy of 78.57%.
Table 3: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on Tiny-ImageNet (Top-1 accuracy)
Model Acc # Params Flops
ViT 55.74 2.69M 174.25M
ViT-LKCA 60.95 (+5.21) 2.69M 170.81M
VIT/0.5M 46.53 0.48M 30.58M
ViT/0.5M-LKCA 50.59 (+4.06) 0.50M 31.34M
VIT/1M 53.46 1.06M 68.36M
ViT/1M-LKCA 57.29 (+3.83) 1.03M 64.82M
VIT/2M 54.58 1.96M 126.09M
ViT/2M-LKCA 60.70 (+6.12) 2.03M 128.66M
VIT/4M 56.16 3.95M 255.47M
ViT/4M-LKCA 62.30 (+6.14) 4.05M 257.43M
VIT/8M 57.39 7.85M 508.72M
ViT/8M-LKCA 63.38 (+5.99) 7.98M 506.00M
Table 3 provides the comparison on Tiny-ImageNet. ViT and ViT-LKCA: The baseline ViT model achieves a Top-1 accuracy of 55.74% with 2.69M parameters and 174.25M Flops. In contrast, the ViT-LKCA model outperforms significantly with a Top-1 accuracy of 60.95% (an increase of 5.21 percentage points) and has the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 46.53% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, incorporating LKCA enhancements, significantly improves accuracy to 50.59% (an increase of 4.06 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): With ViT/4M-LKCA achieving a remarkable 6.14 percentage point improvement and reaching an accuracy of 62.30%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, ViT/8M-LKCA exhibits a significant improvement of 5.99 percentage points, achieving a Top-1 accuracy of 63.38%.
In summary, the LKCA improvement demonstrates significant performance gains across various parameter magnitudes, confirming its effectiveness on ViT models.
Comparison of ViT Variants and LKCA. Next, we compared LKCA with other mainstream variants of Vision Transformer.
First, we provide a comprehensive comparison of LKCA with various mainstream ViT variants on different datasets, including CIFAR-10, CIFAR-100, and SVHN. Each model’s Top-1 accuracy, number of parameters, and Floating Point Operations are presented. In the deep ViT, localized ViT, and lightweight ViT categories, we selected representative models such as Swin-T [53], CaiT-T [69] and MobileViTv2 [70], to compare with standard ViT [3] and our proposed LKCA.
Table 4: Comparison of LKCA with Several Mainstream ViT Variants on Different Datasets (Top-1 accuracy)
Model CIFAR10 CIFAR100 SVHN # Params Flops
ViT-Lite [3] 91.16 70.86 97.53 1.06M 68.36M
Swin-T [53] 91.21 70.37 96.40 1.01M 38.26M
CaiT-T [69] 91.99 70.65 97.53 1.00M 61.16M
MobileViTv2-T [70] 91.46 67.88 96.60 1.00M 7.43M
LKCA-T 92.49 (Ours) 73.40 (Ours) 97.15 (Ours) 1.03M 64.82M
ViT-Small [71] 93.36 72.65 97.90 2.69M 174.25M
Swin-S [53] 93.99 75.34 97.59 2.87M 94.72M
CaiT-S [69] 94.02 74.70 97.79 2.74M 163.55M
MobileViTv2-S [70] 93.48 71.70 96.98 3.92M 24.55M
LKCA-B 94.11 (Ours) 76.50 (Ours) 97.68 (Ours) 2.69M 170.81M
From the information in the table 4. In the context of small-scale models, Swin-T and CaiT-T perform similarly in terms of Top-1 accuracy, while MobileViTv2 lags slightly behind but possesses a significant advantage in computational efficiency. Compared to other small-scale models, LKCA-T achieves a higher Top-1 accuracy. Moving to large-scale models, it is a common trend that they generally outperform their small-scale counterparts. LKCA-B stands out by significantly surpassing ViT and MobileViTv2 in Top-1 accuracy, and it slightly outperforms CaiT-S and Swin-S. Overall, in CIFAR10 and CIFAR100, LKCA demonstrates higher Top-1 accuracy compared to other models. However, there is no clear advantage observed in the SVHN dataset.
Finally, we have validated the performance of various ViT variants and the LKCA model on the Tiny-ImageNet dataset, while also providing details on the models’ parameter counts and computational complexities in Table 5.
In the category of small-scale models, LKCA-T achieves a Top-1 accuracy of 57.29%, demonstrating excellent performance compared to other models. Its performance surpasses that of several common small-scale models. Moving to the medium-scale models, LKCA-B achieves an impressive Top-1 accuracy of 60.95%, once again showcasing outstanding performance. In the realm of large-scale models, LKCA-L attains a Top-1 accuracy of 63.43%, exhibiting a significant advantage in accuracy compared to models of the same scale. Overall, across various scales, LKCA demonstrates outstanding Top-1 accuracy on the Tiny-ImageNet dataset compared to other competitors.
Table 5: Comparison of Various ViT Variants with LKCA on Tiny-ImageNet Dataset (Top-1 accuracy)
Model Tiny-ImageNet # Params Flops
T2T-T [72] 53.92 1.07M 78.23M
RvT-T [73] 50.65 1.10M 57.61M
Swin-T [53] 54.93 1.06M 38.90M
CaiT-T [69] 54.76 1.03M 61.85M
XCiT-T [74] 56.78 0.96M 51.44M
ViT-Lite [3] 53.46 1.11M 69.64M
DeepViT-T [60] 34.64 0.99M 62.96M
RegionViT-T [59] 54.32 0.97M 29.38M
CrossViT-T [75] 47.03 1.04M 57.59M
LKCA-T 57.29 (Ours) 1.07M 66.10M
T2T-S [72] 41.25 2.56M 52.96M
RvT-S [73] 55.51 2.72M 145.09M
Swin-S [53] 58.61 2.93M 95.55M
CaiT-S [69] 59.21 2.77M 164.46M
XCiT-S [74] 60.09 2.81M 157.54M
ViT-Small [3] 55.74 2.76M 176.06M
DeepViT-S [60] 44.45 2.54M 162.85M
Twins_SVT-S [55] 37.13 2.76M 197.00M
RegionViT-S [59] 53.96 2.86M 53.82M
CrossViT-S [75] 52.70 2.40M 126.11M
LKCA-B 60.95 (Ours) 2.76M 172.78M
T2T-B [72] 58.46 13.45M 853.02M
CvT-B [71] 55.88 6.52M 102.56M
MobileViTv2 [70] 58.28 8.17M 189.77M
Twins_SVT-B [55] 49.24 9.04M 308.74M
RegionViT-B [59] 57.83 12.39M 195.02M
LKCA-L 63.43 (Ours) 12.65M 802.99M
4.2 Semantic Segmentation
Experimental Setting. We use MMSegmentation [76] to conduct our experiments on semantic segmentation. All our experiments are carried out on the ADE20K [77] dataset, testing two algorithms: UperNet [78] and FPN [79]. We refrain from using pre-training and instead train the models from scratch on the dataset. Regarding evaluation metrics, we assess the models based on aAcc and mIoU in 10 high-score categories. In the ablation study of the backbone network, we examine various popular variants of ViT and CNN and explore their performance during 0 - 20,000 iterations.
Segment on ADE20K. We conducted experiments on multiple backbone networks on ADE20K
Table 6: Semantic Segmentation of Vanilla ViT and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params Flops
ViT-B 2000 UPerNet 40.90 22.28 0.144G 0.108T
LKCA-B 2000 UPerNet 45.48 24.65 0.102G 0.115T
ViT-B 4000 UPerNet 47.41 28.08 0.144G 0.108T
LKCA-B 4000 UPerNet 52.67 30.00 0.102G 0.115T
ViT-B 8000 UPerNet 53.43 31.73 0.144G 0.108T
LKCA-B 8000 UPerNet 50.79 33.73 0.102G 0.115T
ViT-B 16000 UPerNet 61.55 39.33 0.144G 0.108T
LKCA-B 16000 UPerNet 65.49 40.60 0.102G 0.115T
Table 6 provides performance metrics for the UPerNet algorithm using different backbones (ViT and LKCA), iteration algorithms, and iteration counts (2000, 4000, 8000, 16000). At each iteration count, LKCA-B consistently outperforms ViT-B in terms of aAcc and mIoU, indicating superior performance of LKCA-B in the UPerNet task. As the iteration count increases, both models show an improvement in performance, but at each iteration count, LKCA-B maintains a leading position. Overall, LKCA-B demonstrates superior performance compared to ViT-B at different iteration counts, with higher pixel accuracy and mean Intersection over Union. Additionally, LKCA-B exhibits advantages in terms of parameter count and computational complexity.
Table 7: Semantic Segmentation of Various ViT Variants and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params Flops
Vit-S 10000 UPerNet 55.55 32.92 57.66M 67.25G
ResNet-50 10000 UPerNet 58.12 35.35 64.12M 0.238T
Twins_PCPVT-S 10000 UPerNet 52.56 29.64 53.30M 0.234T
Swin-S 10000 UPerNet 57.18 35.30 51.90M 0.231T
LKCA-S 10000 UPerNet 60.16 37.16 50.58M 65.94G
Vit-S 10000 FPN 56.56 35.84 37.23M 76.14G
Twins_PCPVT-S 10000 FPN 56.74 33.09 33.43M 49.74G
LKCA-S 10000 FPN 62.55 39.03 30.18M 77.92G
Vit-S 20000 UPerNet 58.51 36.61 57.66M 67.25G
ResNet-50 20000 UPerNet 62.76 37.84 64.12M 0.238T
Twins_PCPVT-S 20000 UPerNet 58.91 38.00 53.30M 0.234T
Swin-S 20000 UPerNet 58.00 37.57 51.90M 0.231T
LKCA-S 20000 UPerNet 64.37 40.09 50.58M 65.94G
Vit-S 20000 FPN 63.36 41.64 37.23M 76.14G
Twins_PCPVT-S 20000 FPN 58.51 36.61 33.43M 49.74G
LKCA-S 20000 FPN 67.21 41.81 30.18M 77.92G
Table 7 presents a performance comparison among different backbones (ViT-S, ResNet-50, Swin-S, Twins PCPVT-S, LKCA-S) under various iteration counts and across two different task settings (UPerNet and FPN). In the UPerNet task, LKCA-S consistently outperforms other models at different iteration counts, exhibiting higher aAcc and mIoU. It achieves superior performance compared to other models. Similarly, in the FPN task, LKCA-S demonstrates excellent performance across different iteration counts, showcasing higher aAcc and mIoU, making it competitive against other models. LKCA-S consistently exhibits outstanding performance across different tasks and iteration settings, while maintaining relatively lower computational complexity and parameter counts.
5 Conclusion
In this study, we re-examine the relationship between the attention mechanism in visual transformers and the large kernel convolutional networks, proposing a new visual attention called Large Kernel Convolutional Attention (LKCA). It employs a single large kernel convolution to simplify attention operations. In comparison to the previous approach, Large Kernel Attention (LKA), which uses multiple small kernel convolutions, our method achieves the same effects but is more straightforward and simpler. LKCA combines the advantages of convolutional neural networks and visual transformers, featuring a larger receptive field, locality, and parameter sharing. We explained the superiority of LKCA from the perspectives of both convolution and attention, providing equivalent code implementations for each view. Experimental results demonstrate that LKCA, implemented from both convolutional and attention perspectives, exhibits comparable performance. We extensively experimented with LKCA variants of ViT on multiple benchmark datasets, including CIFAR-10, CIFAR-100, SVHN, Tiny-ImageNet, and ADE20K, for classification and segmentation tasks. The experimental outcomes reveal that LKCA demonstrates competitive performance in visual tasks.